OpenStudy (dls):

Derivative with 3 variables?

OpenStudy (dls):

\[\large x^\frac{2}{3} +y^\frac{2}{3} =a^\frac{2}{3} \]

zepdrix (zepdrix):

A generally represents a CONSTANT. I doubt that's a third variable in your problem. :D Derivative of a constant issss?

OpenStudy (dls):

00000000000000000000000000000000000000000

zepdrix (zepdrix):

You missed a couple zeros, but yes that's correct :o

OpenStudy (dls):

can we cut the powers directly?

OpenStudy (dls):

lol

zepdrix (zepdrix):

Cut them? :o

OpenStudy (dls):

\[\frac{2}{3}x^{\frac{-1}{2}} +\frac{2}{3}y^{\frac{-1}{2}}=0\]

OpenStudy (dls):

\[\sqrt{\frac{3}{2}}x+\sqrt{\frac{3}{2}}y \frac{dy}{dx}=0 \]

OpenStudy (dls):

\[x+y \frac{dx}{dx}=0\]

OpenStudy (dls):

\[\frac{dy}{dx}=\frac{-x}{y}\]

OpenStudy (dls):

:O pro maths!!

OpenStudy (dls):

:")

zepdrix (zepdrix):

\[\large \frac{2}{3}x^{\frac{-1}{2}} +\frac{2}{3}y^{\frac{-1}{2}}\frac{dy}{dx}=0\]This step looked good. I'm not sure what happened after that :O

OpenStudy (dls):

1/2 meant square root and minus meant reciprocal so did that :D

OpenStudy (dls):

oh DAM SORRY

zepdrix (zepdrix):

But the exponent is only being applied to x :O not the constant in front :O Silly DSL, such unreliable internet access, upgrade to cable D:

OpenStudy (dls):

DSL and DLS hha :D

zepdrix (zepdrix):

XD

OpenStudy (dls):

\[\frac{2}{3\sqrt{x}}+\frac{2}{3\sqrt{y}}=0\]

OpenStudy (dls):

MISSING DY/DX!!*

zepdrix (zepdrix):

Yesss :O scary!

OpenStudy (dls):

\[\frac{dy}{dx}=\frac{9 \sqrt{x} \sqrt{y}}{ \sqrt{x} +\sqrt{y}}\]

OpenStudy (dls):

something like this perhaps O.o

zepdrix (zepdrix):

Something like this perhaps? :o \[\large \frac{2}{3\sqrt{x}}+\frac{2}{3\sqrt{y}}\frac{dy}{dx}=0 \qquad \rightarrow \qquad \frac{dy}{dx}=-\frac{\left(\frac{2}{3\sqrt x}\right)}{\left(\frac{2}{3\sqrt y}\right)}\] There is some more simplification possible it looks like though :D

OpenStudy (dls):

I got \[\frac{dy}{dx}=\frac{3xy}{\sqrt{3}(y+x)}\]

zepdrix (zepdrix):

Oh I just noticed, these are suppose to be cube roots, not square roots right? 2/3 - 1 = -1/3

OpenStudy (dls):

oh yes :P

OpenStudy (dls):

\[\large 2(\frac{1}{3 \sqrt[3]{x}}+\frac{1}{3 \sqrt[3]{y}} ) =0\]

OpenStudy (dls):

oh and take the LCM

zepdrix (zepdrix):

I think it might help to solve this if you leave them as fractional exponents maybe :O \[\large \frac{2}{3}x^{-1/3}+\frac{2}{3}y^{-1/3}\frac{dy}{dx}=0\]Multiplying both sides by 3/2 gives us,\[\large x^{-1/3}+y^{-1/3}\frac{dy}{dx}=0\]Subtracting the x term from both sides gives us,\[\large y^{-1/3}\frac{dy}{dx}=-x^{-1/3}\]

OpenStudy (dls):

\[\frac{3 \sqrt[3]{x}+3 \sqrt[3]{y}}{3 \sqrt[3]{x}}dy/dx=0\]

OpenStudy (dls):

yeah same thing^^

OpenStudy (dls):

y/x ?

zepdrix (zepdrix):

\[\large \frac{dy}{dx}=-\frac{y^{1/3}}{x^{1/3}}\]Should get something like this I think. Hmm.

OpenStudy (dls):

\[\large \frac{dy}{dx}=\frac{\sqrt[3]{y}}{\sqrt[3]{x}}\]

OpenStudy (dls):

oooooooo :D

zepdrix (zepdrix):

Yay team.

OpenStudy (dls):

holdon